Step (1) Rules for Guessing: 1. If y(x) = ek *. guess that yp(I) = Acts then find A 2. If g(x) = sin(kx), guess that y(x) = solve number 2 for me pls.
Answer #1

2. The given differential equation is +2y+54 cos (2x) y Te-x To find C.F. Let =ema - then the auxiliary y. #o be a troial sTo find p. Yp: 4 e-x cosaa D. +2D +5 -4e-> cos 2x (0-1)2+2(0-1) +5 1 = 40-e D2-20+1+20-2+5 cos2x = te-x D2 + 4 Cos 2x 4e-2 le- ezize xo Adil 1 + 41 Felix 1 4 Di seo (1 + ) 110 (1-0/41).** = e zix - ezize I · 1 4iD re - zixe e 4i Similarly, e-zise +4Now y(0)=0 gives co x Ya-e- (cz sin 2x + x sin 2x) + 262 Cos 2x + Sin 2 x) e-se + 2xe-x Cosaa Now yCOD=1 gives 1 = 262 2 =>